Suppose the demand and supply of chickens is given by:

QD = 22,500 – 250P

QS = 5,000 + 100P

a. Find the equilibrium price and quantity.
b. Suppose a quota of 7,000 chickens in imposed. What will be the new equilibrium price? What is the loss to consumers? What is the net gain, if any, to producers?
1
1
Expert's answer
2019-06-10T09:31:08-0400

1) At equilibrium Quantity and price Quantity demanded =Quantity supplied

 22,500 – 250P=5000+100P

17500=350P

50=P

QD=QS=22500-250(50) =10000

2) QS=7000, QD=22500-250P

7000=22500-250P

-15500=-250P

P=62

Loss to consumers =change in consumers surplus =1/2*12(7000+10000) =102000

Gain to producers = change in producers surplus

Producer surplus before=1/2(100*10000) =500000

Producers surplus after=1/2*7000(112+42) =539000

Net gain to producers =39000